# The Nuclear Star Cluster of M 74: a fossil record of the very early stages of a star-forming galaxy

### v2

@@ 5 Results
5.1 Stellar populations of the galactic center: the NSC compared with the host galaxy
5.1.1 Average stellar population properties within a PSF aperture
@@ 5.1.2 Additional tests and comparisons
Additionally, we analyzed the ”original” data cube (defined in Sect. 4.1.2 as the prepared data cube just before applying C2D), to assess the robustness of our results. We first compare the results above with the stellar populations obtained from a central PSF aperture in the original data cube (spectrum indicated as a blue solid line in Fig. 4). This contains the integrated light of the NSC and the host galaxy together and, as expected, gives intermediate stellar populations between the NSC and the host galaxy (Table 1). We can also compare the stellar populations in a central PSF aperture with the surrounding region of the galaxy, both extracted from the original data cube. For that, we calculated the average ages, [M/H] and [Mg/Fe] within a much larger central aperture of 100 pc (excluding the central $3\sigma$ PSF to make sure the contamination from the NSC is negligible). For this, we used the stellar-population maps extracted from the original integrated data cube (Appendix A). These stars are younger (5.6 Gyr) and slightly more metal-rich (0.1-0.1 dex), than the stars in the central PSF aperture extracted from the same original data cube (Table 1). However, the [Mg/Fe] abundance is very similar (0.1 dex). The fact that the original data cube, including the light of stars of both the NSC and the host galaxy, shows gradients with older and more metal-poor stars in the very center, is in perfect agreement with the presence of an older and more metal-poor component in the center (i.e. the NSC), which is not present in the surroundings.
@@
To additionally test how robust our C2D method is, we have also compared our results with the ones from a NSC spectrum obtained by subtracting the underlying host-galaxy light from a PSF integrated spectrum (”original”, in Fig. 4). Following Fahrion et al. (2021), the host-galaxy spectrum was obtained from an annulus surrounding the NSC, with inner and outer radii of 0.8 ($\sim 3\sigma$ PSF) and 2 arcsec. Before subtracting the host-galaxy spectrum from the ”original” spectrum, the annulus spectrum needs to be first rescaled to the light of the host galaxy in the same central PSF aperture where the ”original” spectrum was integrated (brighter than the annulus). While Fahrion et al. (2021) performed a dedicated morphological decomposition to calculate the scaling factor, we used our decomposition from C2D. We obtained again very old light-weighted ages ($8.5\pm 1.1$ Gyr), low metallicities ($0.39\pm 0.04$ dex) and close to solar [Mg/Fe] enhancement ($0.01\pm 0.03$ dex). While these results give small differences (taking into account the uncertainties - slightly younger ages and higher metallicities), they are closer to the results from C2D than what was obtained from the original data cube. While it is not surprising that different methods give slightly different results (see also Sect. 6.1), our method is more sophisticated than simply subtracting spectra. Our method leads to a cleaner NSC spectrum than the simple subtraction, and therefore to larger differences between results from integrated and decomposed spectra. The main reason for this is that C2D performs a fitting spaxel by spaxel and spectral pixel by spectral pixel, while when subtracting the annulus spectrum from the PSF spectrum we are forced to take an ”average” spectrum within the annulus, and the scaling factor is also calculated using integrated light. More importantly, these differences are not critical for the interpretation of our results: the NSC appears to be very old and metal poor, and much older and more metal poor than the host galaxy according to the results obtained with all methods.
5.1.3 Star formation histories (SFHs) of the NSC and comparison with the host galaxy
We analyze now the mass fraction assigned to SSP models of different ages during the fitting process, which were used to calculate the mass-weighted averages in Table 1. This allows us to reconstruct SFHs in terms of the mass in stars of different age bins. We extracted SFHs of the NSC and the host galaxy from pPXF fits of the spectra extracted in our $1\sigma$ PSF aperture (Sect. 4.2). We show these SFHs in Fig. 5, respectively in the top and middle panels. In the bottom panel we show, as a reference, the SFH of the full non-decomposed galaxy, extracted from the same PSF aperture in the original MUSE data cube before the C2D decomposition (blue spectrum in Fig 4).
The NSC shows distinct stellar populations at old ages, with the highest mass fractions distributed in ages between 10 and 11 Gyr. The host galaxy shows an old component made of two peaks of different ages (${\sim}10$ and ${\sim}12$ Gyr) and a relatively recent component with ages of about 2 Gyr and younger. Some intermediate-age populations are also present (${\sim}6$ Gyr). The SFH of the full galaxy (bottom panel) shows a combination of all these components. A word of caution is deserved by the large uncertainties, indicated by the shades in each panel of Fig. 5. These shades cover the region between the 16% and 84% percentiles, calculated on the distribution of the 100 Monte Carlo realizations. Since mass fractions are normalized to the total mass in each SFH (in each fitted spectrum), these errors indicate that some mass fraction assigned to the highest peaks could be, in reality, redistributed to minor peaks. Overall, the SFHs suggest that the center of the galaxy had an intense early phase of star formation, between 14 and 8 Gyr ago, when the NSC stars formed. Afterwards, it experienced a long quiescent phase, with a very low star-formation rate. Only relatively recently (later than 3 Gyr ago), a more active star formation was reactivated in M 74. However, while this episode involved an extended region of the disk and was enhanced in spiral arms, as shown in Fig. 6, it did not affect the NSC, as this remained old and metal poor. Sánchez-Blázquez et al. (2014a) found a very similar SFH for M 74, using the STEllar Content via Maximum A Posteriori likelihood (STECKMAP) spectral-fitting code (Ocvirk et al. 2006).
@@ 5.2 Spatially resolved stellar populations of the host galaxy
Spatially resolved properties of the host galaxy are important to interpret the NSC properties in the context of its surroundings. We mapped the mean age, [M/H] and [Mg/Fe] of the host galaxy, fitting the host-galaxy data cube decomposed with C2D (Sect. 4.1.2 and 4.2). The maps are shown in Fig. 6. Both [M/H] and [Mg/Fe] are presented in logarithmic scale and with respect to solar values (zero values correspond to solar values). The age map (top panel) shows a relatively young host galaxy (with most regions younger than ${\sim}4$ Gyr). A pattern following the spiral arms can be identified, with alternative regions dominated either by young (about ${\sim}2$ Gyr-old or younger) or old stars (older than ${\sim}8$ Gyr). The [M/H] map (middle panel) displays a clear negative radial gradient, with a central metal-rich region of about 500 pc size (supersolar metallicities). Outer regions have alternatively higher and lower values. [Mg/Fe] abundances are mostly anti-correlated with [M/H], showing the lowest values (close to solar) in the central metal-rich ${\sim}500$ pc. Hints of the spiral-arm shape are visible in both [M/H] and [Mg/Fe] maps, whose values do not necessarily follow any correlation or anti-correlation with age. However, the youngest (outer) regions in the analyzed data cube show a strong $\alpha$ enhancement in some metal-poor blobs.
@@ 6.1 The nuclear star cluster in M 74 and its formation
Our stellar population analysis has shown an ancient and metal-poor NSC, surrounded by a much younger and more metal-rich region of the host galaxy. As shown by JWST observations (Hoyer et al. 2023b), the NSC lives in a cavity with no gas or dust, while being hosted by an actively star-forming galaxy with a complex structure of gas and dust filaments. The cavity formed due to some processes either keeping gas and dust out of it or wiping them out of the central region (see also Sect. 6.3). While the presence of a massive black hole in M 74 is still under debate, the presence of a low-luminosity AGN was proposed by different authors (e.g., Dong and De Robertis 2006; She et al. 2017). It is also one of the possible explanations for the presence of the cavity (all gas may have been blown via AGN feedback, especially if the black-hole activity was much stronger in the past) and of the mid-infrared emission found at the center of the galaxy with JWST observations (Hoyer et al. 2023b). Similarly, central quiescent regions (of about 100 pc sizes), surrounded by starbursts, as well as nuclear older ages (within radii from few to ${\sim}50$ pc) than the circumnuclear regions, were previously detected in AGN hosts (e.g., Márquez et al. 2025; da Silva et al. 2025).
@@
In any case, the presence of the cavity may have affected the evolution of the NSC, which could not grow via in-situ star formation due to the lack of gas. However, the surrounding region of the NSC, including stars in the region of the cavity, is much younger and more metal-rich. Our decomposition, supported by the analysis of spectra in the original data cube integrated into different apertures (Sect. 5.1.2), suggests that these younger stars belong to the galaxy disk. Our SFH of the center of the host galaxy suggests that the latter had later and relatively recent episodes of star formation. This did not happen in the NSC, whose stars were all formed at the early stages of the evolution of M 74. Its SFH shows that its stars were all formed more than 8 Gyr ago. Regarding the origin of these old and metal-poor stars, we cannot say if they were formed in situ a long time ago, when M 74’s gas was still metal-poor. Old ages and low metallicities (although typically lower than what we find here) in NSCs are traditionally explained by globular cluster inspiral (Neumayer et al. 2020; Fahrion et al. 2021, 2022b, 2022a). While globular cluster migration is thought to be the dominant formation mechanism for NSCs less massive than the one in M 74 and hosted by low-mass galaxies, our results do not provide enough information to rule out (or support) this scenario versus a very early in-situ formation (see also Sect. 6.3). Furthermore, the presence of LMXBs in the NSC suggested by the nuclear X-ray emission (see Sect. 3.1), consistent with the NSC old age, provides additional support for the NSC to be one or more globular clusters that migrated to the galactic center.
@@ 6.2 Is M 74’s nuclear star cluster unique?
The NSC in M 74 shows common stellar population properties with other NSCs in the literature. However, it is important to discuss NSC properties in the context of their host galaxies. Fahrion et al. (2021) analyzed NSCs in a sample of 25 ETGs in the Fornax and Virgo clusters, finding several of them as old as M 74’s NSC and very metal-poor. However, while the sample includes galaxies of similar stellar masses to M 74 (Sect. 3), the NSCs with the most similar stellar populations to M 74’s NSC were lower-mass NSCs ($\lesssim 10^{6}$ M) hosted by less massive, mostly quiescent ETGs ($\lesssim 10^{9}$ M). ETGs of similar masses as M 74 host metal-rich NSCs. Among the seven dwarf ETGs in the Fornax cluster, analyzed by Johnston et al. (2020), NSCs with similar ages to the one in M 74 are much more metal-poor. These NSCs and their dwarf host galaxies are systematically more metal-poor than in the M 74 case. M 74’s NSC age is also similar to three NSCs from Fahrion et al. (2022a), who analyzed a sample of nine star-forming dwarf LTGs. While these galaxies are star-forming and have morphological types more similar to M 74, they have much lower stellar masses (${\sim}10^{7}$ to ${\sim}10^{9}$ M). These host galaxies and their NSCs also show systematically lower metallicities than M 74 and its NSC. Some NSCs hosted by galaxies more similar to M 74 in terms of morphological types and masses were analyzed by Kacharov et al. (2018). However, while some have similar metallicities to M 74’s NSC (see also Walcher et al. 2006), they are all relatively young. While there are other spiral galaxies with NSC ages similar to M 74’s, these galaxies are slightly earlier types than M 74 (Rossa et al. 2006).
While previous studies in the literature show other NSCs with similar stellar-population properties, suggesting that M 74’s NSC is not unique, no NSCs hosted by a relatively young, metal-rich, massive star-forming (late-type) spiral galaxy, have been reported to have such old ages. This suggests that the properties of M 74’s NSC are not common in its host-galaxy context, and this NSC may have had a peculiar formation and evolution history.
@@ 7 Conclusions
We found a NSC average age of about 11 Gyr and a total metallicity [M/H] of about 0.5-0.5 dex. The SFH of the NSC shows that it is made up exclusively of old stellar populations. These results indicate that NSC stars formed during the earliest stages of the evolution of M 74, probably spanning 2-3 distinct episodes. While we cannot track the potential migration of these stars from other regions of the galaxy, and we did not find hints of their accretion from galaxy satellites, our results do not show later contributions with stars younger than ${\sim}8$ Gyr. The NSC appears to have remained inactive, in terms of star formation and chemical evolution, for a long time. Interestingly, this NSC lives in a cavity (with a size of ${\sim}300$ pc) devoid of gas and dust, while the rest of the galaxy shows a complex web of gas and dust filaments, where star formation is currently happening. Stellar-population maps of the host galaxy have shown young ages in the spiral arms, and a central very metal-rich ${\sim}500$ pc region, with lower [Mg/Fe] close to solar values. This indicates that the surroundings of the NSC kept evolving over time, while its evolution stopped. The comparison of spectra extracted in a central PSF aperture, for the NSC and the host galaxy, shows that in contrast to the short NSC evolution, the center of the host galaxy had relatively recent star-formation episodes, in addition to the old populations coeval with the NSC.
Finally, the NSC in M 74 is much more metal poor not only than its host galaxy, but also than other NSCs hosted by ETGs of the same mass as M 74. While there are no measurements for spiral galaxies of the same mass as M 74, its NSC is more metal poor than other NSCs in slightly lower-mass LTGs, with stellar masses between $10^{9.5}$ and $10^{10}$ M. The literature measurements mentioned above might be contaminated by the light of the host galaxy. Our spectro-photometric decomposition method allowed us to obtain a cleaner view of the stellar populations of the NSC, otherwise contaminated by the host galaxy (as shown in Table 1 by the analysis of the ”Original” data cube). While these findings would be traditionally interpreted as a dominant NSC formation through globular cluster migration and merger at the center of the galaxy, they also suggest that low metallicities might emerge from in-situ star formation (at early times) in massive LTGs. With the analysis of a larger sample of massive spiral galaxies, we will find out whether the NSC in M 74 is a unique case or if its stellar populations are common in star-forming galaxies of similar masses, providing additional insights on NSC formation.

### v3

@@ 5 Results
5.1 Stellar populations within a central PSF aperture: average properties and SFHs of the NSC compared with the host galaxy
We present here the average stellar population properties and SFHs, both reconstructed from the weights given to the SSP models by the spectral-fitting approach detailed in Sect. 4.2. All these results were extracted from the three spectra integrated within an aperture of $1\sigma$ PSF (Sect. 4.1.2) and shown in Fig. 4.
@@
From the mass weights obtained in the full spectral fitting process, we also reconstructed the SFHs in terms of the stellar mass assigned to different age bins. Compared to the average properties presented above, SFHs were obtained by averaging only the mass weights given to coeval SSPs. We show these SFHs in Fig. 5, respectively in the top and middle panels. In the bottom panel, we show, as a reference, the SFH of the full non-decomposed galaxy, extracted from the same PSF aperture in the kinematic-corrected original MUSE data cube before the C2D decomposition (blue spectrum in Fig 4). The NSC shows distinct stellar populations at old ages, with the highest mass fractions distributed in ages between 10 and 11 Gyr. The host galaxy shows an old component made of two peaks of different ages (${\sim}10$ and ${\sim}12$ Gyr) and a relatively recent component with ages of about 2 Gyr and younger. Some intermediate-age populations are also present (${\sim}6$ Gyr). The SFH of the full galaxy (bottom panel) shows a combination of all these components. A word of caution is deserved by the large uncertainties, indicated by the shades in each panel of Fig. 5. These shades cover the region between the 16% and 84% percentiles, calculated on the distribution of the 100 Monte Carlo realizations. Since mass fractions are normalized to the total mass in each SFH (in each fitted spectrum), these errors indicate that some mass fraction assigned to the highest peaks could be, in reality, redistributed to minor peaks. Overall, the SFHs suggest that the center of the galaxy had an intense early phase of star formation, between 14 and 8 Gyr ago, when the NSC stars formed. Afterwards, it experienced a long quiescent phase, with a very low star-formation rate. Only relatively recently (later than 3 Gyr ago), a more active star formation was reactivated in M 74. However, while this episode involved an extended region of the disk and was enhanced in spiral arms, as shown in Fig. 6, it did not affect the NSC, as this remained old and metal poor. Sánchez-Blázquez et al. (2014a) found a very similar SFH for M 74, using the STEllar Content via Maximum A Posteriori likelihood (STECKMAP) spectral-fitting code (Ocvirk et al. 2006).
@@ 5.1.1 Additional tests and comparisons
We first compare the results above with the stellar populations obtained from a central PSF aperture in the kinematic-corrected original data cube (spectrum indicated as a blue solid line in Fig. 4). This contains the integrated light of the NSC and the host galaxy together and, as expected, gives intermediate stellar populations between the NSC and the host galaxy (Table 1). We can also compare the stellar populations in a central PSF aperture with the surrounding region of the galaxy, both extracted from the original data cube. For that, we calculated the average ages, [M/H] and [Mg/Fe] within a much larger central aperture of 100 pc (excluding the central $3\sigma$ PSF to make sure the contamination from the NSC is negligible). For this, we used the stellar-population maps extracted from the original integrated data cube (Appendix A). These stars are younger (5.6 Gyr) and slightly more metal-rich (0.1-0.1 dex), than the stars in the central PSF aperture extracted from the same original data cube (Table 1). However, the [Mg/Fe] abundance is very similar (0.1 dex). The fact that the original data cube, including the light of stars of both the NSC and the host galaxy, shows gradients with older and more metal-poor stars in the very center, is in perfect agreement with the presence of an older and more metal-poor component in the center (i.e. the NSC), which is not present in the surroundings.
@@
To additionally test how robust our C2D method is, we have also compared our results with the ones from an NSC spectrum obtained by subtracting the underlying host-galaxy light from a PSF integrated spectrum (”original”, in Fig. 4). Following Fahrion et al. (2021), the host-galaxy spectrum was obtained from an annulus surrounding the NSC, with inner and outer radii of 0.8 ($\sim 3\sigma$ PSF) and 2 arcsec. Before subtracting the host-galaxy spectrum from the ”original” spectrum, the annulus spectrum needs to be first rescaled to the light of the host galaxy in the same central PSF aperture where the ”original” spectrum was integrated (brighter than the annulus). While Fahrion et al. (2021) performed a dedicated morphological decomposition to calculate the scaling factor, we used our decomposition from C2D. We obtained again very old light-weighted ages ($8.5\pm 1.1$ Gyr), low metallicities ($0.39\pm 0.04$ dex) and close to solar [Mg/Fe] enhancement ($0.01\pm 0.03$ dex). While these results give small differences (taking into account the uncertainties - slightly younger ages and higher metallicities), they are closer to the results from C2D than what was obtained from the original data cube. While it is not surprising that different methods give slightly different results (see also Sect. 6.1), our method is more sophisticated than simply subtracting spectra. Our method leads to a cleaner NSC spectrum than the simple subtraction, and therefore to larger differences between results from integrated and decomposed spectra. The main reason for this is that C2D performs a fitting spaxel by spaxel and spectral pixel by spectral pixel, while when subtracting the annulus spectrum from the PSF spectrum we are forced to take an ”average” spectrum within the annulus, and the scaling factor is also calculated using integrated light. More importantly, these differences are not critical for the interpretation of our results: the NSC appears to be very old and metal poor, and much older and more metal poor than the host galaxy according to the results obtained with all methods.
@@ 5.2 Spatially resolved stellar populations of the host galaxy
Spatially resolved properties of the host galaxy are important to interpret the NSC properties in the context of its surroundings. We mapped the mean age, [M/H] and [Mg/Fe] of the host galaxy, fitting the host-galaxy data cube decomposed with C2D (Sect. 4.1.2 and 4.2). The maps are shown in Fig. 6. Both [M/H] and [Mg/Fe] are presented in logarithmic scale and with respect to solar values (zero values correspond to solar values). The age map (top panel) shows a relatively young host galaxy (with most regions younger than ${\sim}4$ Gyr). A pattern following the spiral arms can be identified, with alternative regions dominated either by young (about ${\sim}2$ Gyr-old or younger) or old stars (older than ${\sim}8$ Gyr). The [M/H] map (middle panel) displays a clear negative radial gradient, with a central metal-rich region of about 500 pc size (supersolar metallicities). Outer regions have higher and lower values alternatively. [Mg/Fe] abundances are mostly anti-correlated with [M/H], showing the lowest values (close to solar) in the central metal-rich ${\sim}500$ pc. Hints of the spiral-arm shape are visible in both [M/H] and [Mg/Fe] maps, whose values do not necessarily follow any correlation or anti-correlation with age. However, the youngest (outer) regions in the analyzed data cube show a strong $\alpha$ enhancement in some metal-poor blobs.
@@ 6.1 The nuclear star cluster in M 74 and its formation
Our stellar population analysis has shown an ancient and metal-poor NSC, surrounded by a much younger and more metal-rich region of the host galaxy. As shown by JWST observations (Hoyer et al. 2023b), the NSC lives in a cavity with no gas or dust, while being hosted by an actively star-forming galaxy with a complex structure of gas and dust filaments. The cavity formed due to some processes either keeping gas and dust out of it or wiping them out of the central region (see also Sect. 6.3). While the presence of a massive black hole in M 74 is still under debate, the presence of a low-luminosity AGN was proposed by different authors (e.g., Dong and De Robertis 2006; She et al. 2017). It is also one of the possible explanations for the presence of the cavity (all gas may have been blown via AGN feedback, especially if the black-hole activity was much stronger in the past) and of the mid-infrared emission found at the center of the galaxy with JWST observations (Hoyer et al. 2023b). Similarly, central quiescent regions (of about 100 pc sizes), surrounded by starbursts, as well as nuclear older ages (within radii from few to ${\sim}50$ pc) than the circumnuclear regions, were previously detected in AGN hosts (e.g., Márquez et al. 2025; da Silva et al. 2025). Other, more unlikely explanations for the fact that no gas in an amount sufficient to trigger star formation reached the NSC in the last $\sim$8 Gyr, could be a previous overall scarcity of gas in the galaxy (not happening now since the galaxy is overall gas rich), and the lack of a clear and strong bar structure, which would help gas inflow, combined with dynamical processes making gas inflow inefficient.
@@
In any case, the presence of the cavity may have affected the evolution of the NSC, which could not grow via in-situ star formation due to the lack of gas. However, the surrounding region of the NSC, including stars in the region of the cavity, is much younger and more metal-rich. Our decomposition, supported by the analysis of spectra in the original data cube integrated into different apertures (Sect. 5.1.1), suggests that these younger stars belong to the galaxy disk. Our SFH of the center of the host galaxy suggests that the latter had later and relatively recent episodes of star formation. This did not happen in the NSC, whose stars were all formed at the early stages of the evolution of M 74. Its SFH shows that its stars were all formed more than 8 Gyr ago. Regarding the origin of these old and metal-poor stars, we cannot say if they were formed in situ a long time ago, when M 74’s gas was still metal-poor. Old ages and low metallicities (although typically lower than what we find here) in NSCs are traditionally explained by globular cluster inspiral (Neumayer et al. 2020; Fahrion et al. 2021, 2022b, 2022a). While globular cluster migration is thought to be the dominant formation mechanism for NSCs less massive than the one in M 74 and hosted by low-mass galaxies, our results do not provide enough information to rule out (or support) this scenario versus a very early in-situ formation (see also Sect. 6.3). Furthermore, the presence of LMXBs in the NSC suggested by the nuclear X-ray emission (see Sect. 3.1), consistent with the NSC old age, provides additional support for the NSC to be one or more globular clusters that migrated to the galactic center.
@@ 6.2 Is M 74’s nuclear star cluster unique?
The NSC in M 74 shows common stellar population properties with other NSCs in the literature. However, it is important to discuss NSC properties in the context of their host galaxies. Fahrion et al. (2021) analyzed NSCs in a sample of 25 ETGs in the Fornax and Virgo clusters, finding several of them as old as M 74’s NSC and very metal-poor. However, while the sample includes galaxies of similar stellar masses to M 74 (Sect. 3), the NSCs with the most similar stellar populations to M 74’s NSC were lower-mass NSCs ($\lesssim 10^{6}$ M) hosted by less massive, mostly quiescent ETGs ($\lesssim 10^{9}$ M). ETGs of similar masses as M 74 host metal-rich NSCs. Among the seven dwarf ETGs in the Fornax cluster, analyzed by Johnston et al. (2020), NSCs with similar ages to the one in M 74 are much more metal-poor. These NSCs and their dwarf host galaxies are systematically more metal-poor than in the M 74 case. M 74’s NSC age is also similar to that of three NSCs from Fahrion et al. (2022a), who analyzed a sample of nine star-forming dwarf LTGs. While these galaxies are star-forming and have morphological types more similar to M 74, they have much lower stellar masses (${\sim}10^{7}$ to ${\sim}10^{9}$ M). These host galaxies and their NSCs also show systematically lower metallicities than M 74 and its NSC. Some NSCs hosted by galaxies more similar to M 74 in terms of morphological types and masses were analyzed by Kacharov et al. (2018). However, while some have similar metallicities to M 74’s NSC (see also Walcher et al. 2006), they are all relatively young. While there are other spiral galaxies with NSC ages similar to M 74’s, these galaxies are slightly earlier types than M 74 (Rossa et al. 2006).
While previous studies in the literature show other NSCs with similar stellar-population properties, suggesting that M 74’s NSC is not unique, no NSCs hosted by a relatively young, metal-rich, massive star-forming (late-type) spiral galaxy, have been reported to have such old ages. This suggests that the properties of M 74’s NSC are not common in its host-galaxy context, and this NSC may have had a peculiar formation and evolution history. It is quite extraordinary that in such a gas-rich spiral galaxy, no significant amount of gas triggered any in-situ star formation in the NSC over the last $\sim$8 Gyr. The inefficiency of gas inflow to the nucleus in this gas-rich star-forming galaxy is what makes M 74 unusual, resulting in its ancient NSC.
@@ 7 Conclusions
We found an NSC average age of about 11 Gyr and a total metallicity [M/H] of about 0.5-0.5 dex. The SFH of the NSC shows that it is made up exclusively of old stellar populations. These results indicate that NSC stars formed during the earliest stages of the evolution of M 74, probably spanning 2-3 distinct episodes. While we cannot track the potential migration of these stars from other regions of the galaxy, and we did not find hints of their accretion from galaxy satellites, our results do not show later contributions with stars younger than ${\sim}8$ Gyr. The NSC appears to have remained inactive, in terms of star formation and chemical evolution, for a long time. Interestingly, this NSC lives in a cavity (with a size of ${\sim}300$ pc) devoid of gas and dust, while the rest of the galaxy shows a complex web of gas and dust filaments, where star formation is currently happening. Stellar-population maps of the host galaxy have shown young ages in the spiral arms, and a central very metal-rich ${\sim}500$ pc region, with lower [Mg/Fe] close to solar values. This indicates that the surroundings of the NSC kept evolving over time, while its evolution stopped. The comparison of spectra extracted in a central PSF aperture, for the NSC and the host galaxy, shows that in contrast to the short NSC evolution, the center of the host galaxy had relatively recent star-formation episodes, in addition to the old populations coeval with the NSC.
Finally, the NSC in M 74 is much more metal-poor not only than its host galaxy, but also than other NSCs hosted by ETGs of the same mass as M 74. While there are no measurements for spiral galaxies of the same mass as M 74, its NSC is more metal-poor than other NSCs in slightly lower-mass LTGs, with stellar masses between $10^{9.5}$ and $10^{10}$ M. The literature measurements mentioned above might be contaminated by the light of the host galaxy. Our spectro-photometric decomposition method allowed us to obtain a cleaner view of the stellar populations of the NSC, otherwise contaminated by the host galaxy (as shown in Table 1 by the analysis of the ”Original” data cube). While these findings would be traditionally interpreted as a dominant NSC formation through globular cluster migration and merger at the center of the galaxy, they also suggest that low metallicities might emerge from in-situ star formation (at early times) in massive LTGs. With the analysis of a larger sample of massive spiral galaxies, we will find out whether the NSC in M 74 is a unique case or if its stellar populations are common in star-forming galaxies of similar masses, providing additional insights on NSC formation.